One number is 10 more than the other number. Twice the smaller number is 13 more than the larger number. What are the numbers?
please help I need it by today thank you

Answer by rapaljer(4671)

For additional help in Word Problems, see my website by clicking on my tutor name anywhere in algebra.com, look for Basic Algebra, then look for Samples of Basic Algebra: One Step at a Time, then look in Chapter 1 for the appropriate topics. You will find a complete explanation, with examples, exercises, and answers to ALL exercises!!

Okay, here we go:

Let x = the smaller number
x+10 = larger number

"Twice smaller IS larger number plus 13 more."
2x = x+10 + 13
2x = x + 23

Subtract x from each side:
2x -x = x-x +23
x= 23= Smaller number
x+10 = 33= Larger number

Check: Twice the smaller number is 46 which is 13 more than 33. It checks.

R^2 at SCC